# The Peekaboo galaxy: new SALT spectroscopy and implications of archive HST data

### v1

@@ Appendix D Line intensities and element abundances in the E and W H ii regions of Peekaboo dwarf
The obtained abundance ratios of N/O, S/O, Ne/O, Ar/O in all three subsystems are consistent to each other. We can calculate their weighted averages as the characteristic values of the whole galaxy Peekaboo. The results are as follows: log(N/O)= –1.450.03, log(Ne/O)= –0.870.05; log(S/O) = -1.720.07, log(Ar/O) = –2.330.09. These averages are well consistent with the abundances ratios for the lowest metallicities, compiled by Izotov et al. (2006, in their Figure 10, left panel).
@@ Appendix E HST data: identification of the XMP massive hot stars and supergiants
We plot in Fig. 8 positions of all HST stars within the Peekaboo body in the diagram MV versus colour . We mark by their numbers from Table 4 all ’blue’ hot stars and the four supergiants with M 6.0 mag. Here, for both MV and , only extinction in the Galaxy is accounted for, corresponding to E(B-V) = 0.062 (Schlafly, Finkbeiner 2011), and the related AV = 0.186 mag and E() = 0.066 mag. As one can see in the bottom of Table 2, C(H) for all three systems is the same within rather small uncertainties. We adopt its average value of 0.17. Then, for stars residing within the borders of these H ii regions, one can expect the elevated extinction and reddening, in addition to the adopted Milky Way values, AV = 0.16 mag and E()=0.057 mag. In fact, we do not know how this extinction is distributed within the E and W regions. If it is localised within the ’blue nebulosities’, then this ’reddening’ correction is probably applicable only to stars residing close to these ’blue nebulosities’.

### v2

@@ Appendix D Line intensities and element abundances in the E and W H ii regions of the Peekaboo dwarf
The obtained abundance ratios of N/O, S/O, Ne/O, Ar/O in all three subsystems are consistent to each other. We can calculate their weighted averages as the characteristic values of the whole Peekaboo galaxy. The results are as follows: log(N/O)= –1.450.03, log(Ne/O)= –0.870.05; log(S/O) = -1.720.07, log(Ar/O) = –2.330.09. These averages are well consistent with the abundances ratios for the lowest metallicities, compiled by Izotov et al. (2006, in their Figure 10, left panel).
@@ Appendix E HST data: Identification of the XMP massive hot stars and supergiants
We plot in Fig. 8 positions of all HST stars within the Peekaboo body in the diagram MV versus colour . We mark by their numbers from Table 4 all ’blue’ hot stars and the four supergiants with M –6.0 mag. Here, for both MV and , only extinction in the Galaxy is accounted for, corresponding to E(B-V) = 0.062 (Schlafly, Finkbeiner 2011), and the related AV = 0.186 mag and E() = 0.066 mag. As one can see in the bottom of Table 2, C(H) for all three systems is the same within rather small uncertainties. We adopt its average value of 0.17. Then, for stars residing within the borders of these H ii regions, one can expect the elevated extinction and reddening, in addition to the adopted Milky Way values, AV = 0.16 mag and E()=0.057 mag. In fact, we do not know how this extinction is distributed within the E and W regions. If it is localised within the ’blue nebulosities’, then this ’reddening’ correction is probably applicable only to stars residing close to these ’blue nebulosities’.